OpenStudy (anonymous):

Find the x-intercept, y-intercept, and the slope of the line 2x + y = 2. Sketch the graph of 2x + y = 2.

OpenStudy (aroub):

if u want to find the y-intercept u make the x=0 and when u want to find the x-intercept u make y=0

OpenStudy (anonymous):

And the slope is the coefficient of x when put in slope-intercept form.

OpenStudy (anonymous):

so 2(0) + y =2 ?

OpenStudy (aroub):

yes :)

OpenStudy (anonymous):

0 + y = 2 y= 2 ??

OpenStudy (anonymous):

That's right. Now do the same for the x-intercept: 2x+0=2

OpenStudy (anonymous):

are the y intercepts (0,2) and (2,0) then?

OpenStudy (anonymous):

Almost. The y-intercept is (0,2). But the x-intercept is (1,0).

OpenStudy (aroub):

yes :)

OpenStudy (anonymous):

to find the equation for the x intercepts do you replace the x with zero?

OpenStudy (anonymous):

0 + y = 2

OpenStudy (aroub):

to find the x-intercept u replace the y with zero

OpenStudy (anonymous):

y-intercept is replacing the x with 0 x-intercept is replacing the y with 0